# LOLA – An Open-Source Massively Multilingual Large Language Model

### v4

@@ Abstract
Nikit Srivastava, Denis Kuchelev, Tatiana Moteu Ngoli, Kshitij Shetty, Michael Röder, Hamada M. Zahera, Diego Moussallem, Axel-Cyrille Ngonga Ngomo Data Science Group, Paderborn University, Germany <firstname.lastname>@upb.de
@@ 1 Introduction
Large Language Models (LLMs) have shown tremendous capability across a diverse set of tasks in recent years Radford et al. (2019); Kaddour et al. (2023). This progress has propelled research, with many chat-based LLMs 111 ChatGPT: chat.openai.com; LLAMA: llama.meta.com; Mistral: mistral.ai; Gemini: gemini.google.com; Deepseek: deepseek.com. gaining popularity among general users. However, concerns remain, particularly regarding their accessibility for multilingual usage Joshi et al. (2020) and open-source licensing policies Liu et al. (2023). The number of competent LLMs significantly decreases for languages other than English Üstün et al. (2024). This, combined with the curse of multilinguality Conneau et al. (2020), means non-English speakers often have access to inferior systems. Additionally, many new models Jiang et al. (2023); Achiam et al. (2024); Dubey et al. (2024) are pay-to-use, require personal information, or do not fully disclose training details, creating significant hurdles for multilingual research.
To advance multilingual language modeling, we introduce LOLA 222 Source Code: github.com/dice-group/LOLA; Model Weights: huggingface.co/dice-research/lola_v1. , a massively multilingual model that follows a GPT-style Radford et al. (2019) decoder-only architecture with sparse Mixture-of-Experts (MoE)layers Shazeer et al. (2017). MoE architectures have shown strong performance on learning the underlying structure of the data  Chen et al. (2022) but their application in multilingual LLMs remains underexplored. MoE models can effectively increase model capacity with minimal additional computational cost, offering the possibility of leveraging implicit clusters like language family groups and playing a crucial role in addressing the challenges of multilinguality.
Language family groups, consisting of languages sharing common ancestral roots, offer opportunities for enhancing language models. Despite linguistic diversity, these families exhibit structural, syntactic, and semantic similarities Rowe and Levine (2015) that can be exploited to improve performance across related languages.
While modern LLMs Devlin et al. (2019); Liu et al. (2019); Radford et al. (2019); Raffel et al. (2023) should learn effectively from multilingual data, Conneau et al. (2020) discovered that as the number of languages increased, models struggled to generalize unless their capacity was significantly expanded. This phenomenon was dubbed the curse of multilinguality. To address this challenge, we employ the Mixture-of-Experts (MoE)architecture in LOLA. Our goal is to leverage MoE’s strengths to exploit the phylogenetic structure of languages and achieve better prediction performance. In particular, the shared and non-shared parameters of MoE-based models offer a promising approach to mitigating the curse of multilinguality by increasing capacity while remaining compute efficient Shazeer et al. (2017). By exploiting language families, we aim to close gaps in current models—particularly for low-resource languages—by enhancing cross-linguistic transfer learning.
@@ 3.1 Routing Mechanism in MoE Layers
where $h(x)$ contains the logit vectors for all experts, $W_{g}$ is the gating weight matrix, and $x$ is the input. The logit vector and gating probability of the $i$-th expert is denoted by $h(x)_{i}$ $G_{i}(x)$ respectively.
@@ 3.2 Training and Loss Functions
$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\cdot\sum_{t=1}^{T}\log p(s_{t}\mid s_{<t}),$ (8)
@@ 9 Acknowledgements
This work has been supported by the Ministry of Culture and Science of North Rhine-Westphalia (MKW NRW) within the project SAIL under the grant no NW21-059D; and German Federal Ministry of Education and Research (BMBF) within the EuroStars project E! 114154 PORQUE under the grant no 01QE2056C. The authors gratefully acknowledge the computing time provided to them on the high-performance computer noctua2 Bauer et al. (2024) at the NHR Center PC2. These are funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for the national high-performance computing at universities (www.nhr-verein.de/unsere-partner). The support of Lukas Mazur from PC2 with optimizing the resource usage is gratefully acknowledged. Finally, we extend our sincere gratitude to Dr. Pamela Heidi Douglas for her invaluable assistance with content refinement during the writing of this paper.

### v5

@@ Abstract
Nikit Srivastava1, Denis Kuchelev, Tatiana Moteu Ngoli1, Kshitij Shetty2, Michael Röder1, Hamada M. Zahera1, Diego Moussallem1, Axel-Cyrille Ngonga Ngomo1 Data Science Group, Paderborn University, Germany 1 {nikit.srivastava, tatiana.moteu, michael.roeder, hamada.zahera, diego.moussallem, axel.ngonga}@upb.de 2 kshitij@mail.upb.de
@@ 1 Introduction
Large Language Models (LLMs) have shown tremendous capability across a diverse set of tasks in recent years Radford et al. (2019); Kaddour et al. (2023). This progress has propelled research, with many chat-based LLMs 111 ChatGPT: chat.openai.com; LLAMA: llama.meta.com; Mistral: mistral.ai; Gemini: gemini.google.com; Deepseek: deepseek.com. gaining popularity among general users. However, concerns remain, particularly regarding their accessibility for multilingual usage Joshi et al. (2020) and open-source licensing policies Liu et al. (2023). The number of competent LLMs significantly decreases for languages other than English Üstün et al. (2024). This, combined with the curse of multilinguality—a phenomenon in which the ability of models to generalize across multiple languages diminishes unless their capacity is significantly expanded Conneau et al. (2020)—means non-English speakers often have access to inferior systems. Additionally, many new models Jiang et al. (2023); Achiam et al. (2024); Dubey et al. (2024) are pay-to-use, require personal information, or do not fully disclose training details, creating significant hurdles for multilingual research.
To advance multilingual language modeling, we introduce LOLA,222 Source Code: github.com/dice-group/LOLA; Model Weights: huggingface.co/dice-research/lola_v1. a massively multilingual model that follows a GPT-style Radford et al. (2019) decoder-only architecture with sparse Mixture-of-Experts (MoE)layers Shazeer et al. (2017). MoE architectures have shown strong performance on learning the underlying structure of the data  Chen et al. (2022) but their application in multilingual LLMs remains underexplored. MoE models can effectively increase model capacity with minimal additional computational cost, offering the possibility of leveraging implicit clusters like language family groups and playing a crucial role in addressing the challenges of multilinguality.
Language family groups, consisting of languages sharing common ancestral roots, offer opportunities for enhancing language models. Despite linguistic diversity, these families exhibit structural, syntactic, and semantic similarities Rowe and Levine (2015) that can be exploited to improve performance across related languages. Our goal is to leverage MoE’s strengths to exploit the phylogenetic structure of languages and achieve better prediction performance. In particular, the shared and non-shared parameters of MoE-based models offer a promising approach to mitigating the curse of multilinguality by increasing capacity while remaining compute efficient Shazeer et al. (2017). By exploiting language families, we aim to close gaps in current models—particularly for low-resource languages—by enhancing cross-linguistic transfer learning.
@@ 3.1 Routing Mechanism in MoE Layers
where $h(x)$ contains the logit vectors for all experts, $W_{g}$ is the gating weight matrix, and $x$ is the input. The logit vector and gating probability of the $i$-th expert is denoted by $h(x)_{i}$ and $G_{i}(x)$ respectively.
@@ 3.2 Training and Loss Functions
$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\cdot\sum_{t=1}^{T}\log p(s_{t}\mid s_{<t}).$ (8)
@@ 9 Acknowledgements
This work has been supported by the Ministry of Culture and Science of North Rhine-Westphalia (MKW NRW) through the project SAIL (grant no. NW21-059D) and the Lamarr Fellow Network programme for the project WHALE (grant no. LFN 1-04). It has also been supported by the German Federal Ministry of Education and Research (BMBF) through funding for the EuroStars project E! 114154 PORQUE (grant no. 01QE2056C). The authors gratefully acknowledge the computing time provided to them on the high-performance computer noctua2 Bauer et al. (2024) at the NHR Center PC2. These are funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for the national high-performance computing at universities (www.nhr-verein.de/unsere-partner). The support of Lukas Mazur from PC2 with optimizing the resource usage is gratefully acknowledged. Finally, we extend our sincere gratitude to Dr. Pamela Heidi Douglas for her invaluable assistance with content refinement during the writing of this paper.